OpenStudy (anonymous):

A cone of radius 3 cm and slant height 6cm is cut into four identical pieces. Calculate the total surface area of the four pieces.

OpenStudy (anonymous):

Total surface area of the pieces = surface area of cone..

OpenStudy (anonymous):

base area = PI *r*r surface area = PI*r*s PI = 22/7 r= 3 cm S = 6 cm Total area = base area + surface area area of each piece = total area / 4

OpenStudy (anonymous):

I forgot to add the inside area add inside area also for each piece = 1/2 * base* height

OpenStudy (anonymous):

Is it a solid cone....?

OpenStudy (anonymous):

Its not written in the question

OpenStudy (anonymous):

When it's a solid cone it's surface area = area of four pieces